# Divergence-Kernel method for linear responses and diffusion models

### v1

@@ Abstract.
We derive the divergence-kernel formula for the linear response (parameter-derivative of marginal or stationary distributions) of random dynamical systems, and formally pass to the continuous-time limit. Our formula works for multiplicative and parameterized noise over any period of time; it does not require hyperbolicity. Then we derive a pathwise Monte-Carlo algorithm for linear responses. With this, we propose a forward-only diffusion generative model and test on simple problems.
@@
Keywords. Linear response, Divergence method, Likelihood ratio, SDE, Diffusion model.
@@ 1.1. Main results
This paper rigorously derives the divergence-kernel formula for the linear response of discrete-time random dynamical systems. Then we formally pass to the continuous-time limit. Our result gives a fundamental tool for the optimization of diffusion processes; in particular, from now on, we can use the gradient-based method to optimize marginal or stationary densities with respect to the diffusion coefficients. In particular, we propose a simple framework for diffusion generative models. Our main theoretical result is
{theorem}
[formal divergence-kernel formula for linear responses of SDEs] Consider the SDE in $\mathbb{R}^{M}$ parameterized by $\gamma$,
@@
$\begin{split}d\nu=\left((\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha)\nu-\nabla\textnormal{div}F+\nabla^{2}\sigma\nabla\sigma+\nabla\sigma\Delta\sigma\right)dt-(\nabla\sigma\nu^{T}+\nabla^{2}\sigma+\alpha\sigma^{-1})dB.\end{split}$
with initial condition $\nu_{0}=\nabla\log h_{0}(x_{0})$, where $h_{t}$ is the probability density of $x_{t}$. Then $\nabla\log h_{t}(x_{t})=\mathbb{E}\left[\nu_{t}\middle|x_{t}\right]$, and
@@
Here $\Delta\sigma$ is the Laplacian of $\sigma$, $B$ is the Brownian motion, and the SDE is Ito. The time integrations are the limits of Equations˜9 and 10. More generally, $\nu$ is an estimator for the score of the SDE at $\gamma=0$, which can be replaced by other divergence-kernel expressions from [24].
The main advantages of this result are:
@@
    It computes derivative of the marginal density, not an averaged observable;
@@
    The expression in pathwise, so we can compute by Monte-Carlo type algorithms in high-dimensions;
@@
    The diffusion coefficient $\sigma$ can depend on the parameter $\gamma$ and the location $x$;
@@
    The size of the estimator does not grow exponentially to $T$, so we can easily extend our result to stationary distributions (see ˜Remark);
@@
    It does not require hyperbolicity (in the sense of dynamical systems).
As a less obvious application, we use this result to construct a new diffusion generative model, where we directly optimize the marginal distribution to reduce its KL-divergence with respect to the data. Importantly, we can optimize the diffusion coefficients of multiplicative noise.
@@ 1.2. Literature review
The path-perturbation method computes the path-perturbation under fixed randomness (such as Brownian motion), then averages over many realizations of randomness. This is also known as the ensemble method or the stochastic gradient method [7, 15]. It also includes the backpropagation method, which is the basic algorithm for machine learning. However, it is expensive for chaotic or unstable system; the work-around is to artificially reduce the size of the path-perturbation, such as shadowing or clipping methods [26, 29, 30, 8], but they all introduce systematic errors. Moreover, it computes the derivative of an averaged observable, but does not give the derivative of the probability density.
The divergence method is also known as the transfer operator method, since the perturbation of the measure transfer operator is some divergence [10, 36, 38]. However, it is expensive for high-dimensional systems with contracting directions.
The kernel-differentiation method works only for random systems. It is also known as the Cameron-Martin-Girsanov theorem [4, 17] or the likelihood ratio method [33, 31, 11]. However, it cannot handle multiplicative noise or perturbation on the diffusion coefficients. It is also expensive when the noise is small. We gave an ergodic version and a foliated version in [25].
@@
Our path-kernel method in [23] gives the pathwise linear response formula of the diffusion coefficients. It is good at systems with not too small noise and not too much unstableness, it does not require hyperbolicity. However, it can be expensive when the noise is small and unstableness is big. Moreover, only part of the expression is derivative on the terminal densities. The path-kernel idea was also used in the Bismut-Elworthy-Li formula for the derivative with respect to the initial conditions [3, 6, 32].
@@
Our divergence-kernel method in [24] gives the score (derivative of marginal density) of random dynamical systems and SDEs. It works for non-hyperbolic and multiplicative noise systems. It is good at systems with not too much contraction and not too little noise. However, it can be expensive when the noise is small and the contraction is strong. We use this result for the forward-backward diffusion models via nonlinear SDEs [18]. This paper extends this result to linear responses.
Our new linear response formula opens up the possibilities of redesigning the current framework of diffusion models. The current structure of the diffusion model involves a forward process and a backward process [35, 16]. Part of the reason is that previously we did not have a gradient tool to optimize marginal distributions with respect to the diffusion coefficient or under multiplicative noise. If we can directly compute such a gradient, we can directly optimize the diffusion process so that the marginal distribution matches the data. We will demonstrate this idea on some simple problems.
@@ 1.3. Structure of paper
Section˜5 proposes a diffusion model which involves only two forward processes. It uses the divergence-kernel formula (not the backpropagation method, which belongs to the path-perturbation method) to perform a gradient descent on the parameters of an SDE. Its loss function is the KL-divergence between the data and the marginal distribution of the SDE. The logic is more straightforward than the current diffusion models, and the memory cost is less.
@@ 2. Notations and assumptions
where $\left|g_{*}\right|$ is the determinant of the matrix.
@@
The density of the initial distribution $h_{0}$, the probability kernel $k$, drift $f$, and diffusion $\sigma$ are $C^{1}$ jointly in $\gamma$ and $x$.
Our derivation is rigorous for the discrete-time case, but is formal when passing to the continuous-time limit, so we do not list the technical assumptions here, and leave the rigorous proof to a later paper. We just assume that all integrations, averages, and change of limits are legit. We still organize our continuous-time results by lemmas and theorems, but this is just a style to organize the paper and does not imply rigorousness.
@@ 3.1. Divergence formula One-step system
That being said, we can certainly mix the kernel and divergence linear response formulas in one-step, then repeatedly apply to many-step systems, and then pass to the continuous-time limit. That approach might be beneficial if we later work with anisotropic noise, where $\sigma$ is a matrix, and each parameter controls only a low-dimensional structure, so we do not have the $M$ factor. Another scenario where this approach might be more useful is that the noise is low-dimensional. Bot situations incur other difficulties, so we will defer to later papers.
@@ 3.2. Discrete-time divergence-kernel formula for linear responses
$\begin{split}\nu_{0}=\nabla\log h_{0}(x_{0}),\quad\textnormal{}\quad\nu_{n+1}=(1-\alpha_{n+1})g_{b_{n}}^{*-1}\left(\nu_{n}-\textnormal{div}g_{b_{n}*}(x_{n})\right)+\frac{\alpha_{n+1}\nabla\log k(b_{n})}{\sigma_{n}(x_{n})}.\end{split}$
@@
Note that we need the score of $h_{0}$, or $\nu_{0}$. Only for large $T$ and systems with fast decorrelation, we can use any $\nu_{0}$ and the impact would be negligible. See ˜Remark for more details.
@@ 3.3. Preparations for SDEs
Ignoring terms that go to zero in the continuous-time limit, [24] derived explicit expressions of some terms for time-discretized SDEs,
@@
{lemma}
[formal divergence-kernel fomula for score of SDEs [24]] For the SDE in $\mathbb{R}^{M}$,
@@ 3.4. Continuous-time divergence-kernel formula for linear response
Then substitute to prove the corollary. ∎
{lemma}
[expression of $\delta x^{\gamma}_{n}$] Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$,
@@
With these, we can formally pass Definition˜1 to the continuous-time limit.
See 1.1
Proof.
Substitute Sections˜3.4, 3.4 and 3.4 into Definition˜1, we get
@@
We also formally pass to the infinite-time limit of stationary distributions $h^{\gamma}$. We need to further assume that $F_{t}$ and $\sigma_{t}$ are stationary, and that there is $T^{\prime}$ such that $\alpha_{t}$ only depends on $x_{t}\sim x_{T^{\prime}}$, so $(\alpha,\nu,x)$ are jointly stationary. {corollary} [formal ergodic divergence-kernel linear response] For any $C^{2}$ observable function $\Phi:\mathbb{R}^{M}\rightarrow R$, let $h^{\gamma}$ be the stationary measure under parameter $\gamma$, then the linear response of the averaged observable can be expressed almost surely on one path:
@@ 3.5. Discussions
4. Numerical examples
@@
$\begin{split}dx=\beta_{t}(\gamma-x)dt+\sqrt{\beta_{t}}(0.5+\exp\left((\gamma-x)^{2}\right)dB,\quad\textnormal{where}\quad\beta_{t}:=1+3t.\end{split}$
The initial distribution is $x_{0}\sim N(0,1)$. Set $T=0.1$, ${\Delta t}=0.01$, $\alpha=10$. Note that the problem is symmetric for $\gamma$.
We are only interested in terminal conditions $x_{T}$ in the interval [2,2][-2,2], which is further partitioned into $10$ subintervals. We simulate many sample paths, then count the number of paths whose $x_{T}$ fall into a specific subinterval, to obtain the empirical distribution of $x_{T}$ on subintervals. Then we compute the linear response $\delta\log h_{T}$ by averaging the divergence-kernel formula over these paths, and see if our results reflect the trend of $\log h_{T}$ from the empirical distribution. We also use the average of $\nu_{T}$ to compute the scores $\nabla\log h_{T}$. Each subinterval contains at least 180 samples. The $\delta\log h_{T}$ and $\nabla\log h_{T}$ computed for different $\gamma$ and $x$’s are shown in Figure˜1. Both correctly reflect the trend of $\log h_{T}$.
Figure 1. Divergence-kernel method for the linear responses and scores of the 1-dimensional SDE with multiplicative noise, $T=0.3$. The dots are $\log h_{T}$. Each short line is a linear response or score computed by the divergence-kernel algorithm, averaged on paths whose terminal conditions $x_{T}$ fall into the same subinterval. Left: linear responses $\delta\log h_{T}$. Right: scores $\nabla\log h_{T}$ .
@@
We use ˜Remark to compute the linear response of the stationary density of the Lorenz 96 model [14]. The dimension of the system is $M=40$. The SDE is
$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+8+\gamma-0.01(x^{i})^{2}\right)dt+0.5+\exp\left((x-\gamma)^{2}\right)dB^{i},$
$\textnormal{where}\quad i=1,\ldots,M;\quad x_{0}\sim{\mathcal{N}}(0,I).$
Here $i$ labels different directions in $\mathbb{R}^{M}$, and it is assumed that $x^{-1}=x^{M-1},x^{0}=x^{M}$ and $x^{M+1}=x^{1}$. We added noise and the $0.01(x^{i})^{2}$ term, which prevents the noise from carrying us to infinitely far away. We choose the observable function to be
@@
We use the Euler integration scheme with ${\Delta t}=0.002$. A typical orbit is in Figure˜2. Then we use the ergodic linear response formula ˜Remark to compute the linear response of the observable averaged on the stationary measure. The averaged observable $\Phi_{avg}$ and the linear response $\delta\int\Phi h^{\gamma}$ are averaged over 7 orbits of length $T=400$. The decorrelation time is $W=1.5$. The results shown in Figure˜2; the linear response computed using the divergence-kernel method is correct.
@@
5. A simple forward-only diffusion model
We consider the problem in which we are given $N_{data}$ many samples $y_{k}\in\mathbb{R}^{M}$, whose true distribution is unknown. Then we use gradient descent to find an SDE such that it maps the Gaussian distribution to match the data.
First, we describe how to generate samples. Note that the information in this paragraph will not be available when we infer the SDE. The samples are generated by the SDE

### v2

# Divergence-kernel method for linear responses of densities and generative models

@@ Abstract.
We derive the divergence-kernel formula for the linear response of random dynamical systems. Specifically, the pathwise expression is for the parameter-derivative of the marginal or stationary density, not an averaged observable. Our formula works for multiplicative and parameterized noise over any period of time; it does not require hyperbolicity. Then we derive a Monte-Carlo algorithm for linear responses.
We develop a new framework of generative models, DK-SDE, where the model is a parameterized SDE, that (1) directly uses the KL divergence between the empirical data distribution and the marginal density of the SDE as the training objective, and (2) accommodates parametrizations in both drift and diffusion over a long time span, allowing prior structural knowledge to be incorporated explicitly. The optimization is done by gradient-descent enabled by the divergence-kernel method, which involves only forward processes and therefore substantially reduces memory cost. We demonstrate the new model on a 20-dimensional Lorenz system.
@@
Keywords. Linear response, Divergence method, Likelihood ratio, SDE, Generative model.
@@ 1.1. Main results
This paper rigorously derives the divergence-kernel formula for the linear response in discrete-time random dynamical systems. It is a pathwise expression of the derivative of marginal or stationary density with respect to parameters in the drift or diffusion coefficients. Then we formally pass to the continuous-time limit in Section˜1.1.
[…]
Consider the SDE in $\mathbb{R}^{M}$ parameterized by $\gamma$,
@@
$\begin{split}d\nu=\left((\nabla\sigma\nabla\sigma^{T}-\nabla F^{T}-\alpha)\nu-\nabla\textnormal{div}F+\nabla^{2}\sigma\nabla\sigma+\nabla\sigma\Delta\sigma\right)dt-(\nabla\sigma\nu^{T}+\nabla^{2}\sigma+\alpha\sigma^{-1})dB\end{split}$
with initial condition $\nu_{0}=\nabla\log h_{0}(x_{0})$, where $h_{t}$ is the probability density of $x_{t}$. Then $\nabla\log h_{t}(x_{t})=\mathbb{E}\left[\nu_{t}\middle|x_{t}\right]$. Moreover, let $\delta(\cdot):=\partial(\cdot)/\partial\gamma$, then
@@
Here $\Delta\sigma$ is the Laplacian of $\sigma$, $B$ is the Brownian motion, and the SDE is Ito. In this paper $\sigma$ is assumed to be a scalar function uniformly bounded away from 0, and extension to matrices is conceptually straightforward. The time integrations are the limits of Equations˜9 and 10. For large $T$, we can just pick $\alpha$ larger than the magnitude of the strongest contraction rate, which stabilizes $\nu$ from growing exponentially fast. In fact, for memory-light implementations we restrict to $\alpha_{t}=\alpha(x_{t})$ or constant $\alpha$; the backward-adapted freedom is mainly a theoretical gauge. The conditional expectation is taken over all paths ending at the same point; it turns pathwise quantities into terminal density derivative. The main advantages of this formula are:
@@
    Compared with the path-kernel method in [30, 29], it computes derivative of the marginal density, not an averaged observable. Moreover, it involves forward processes so the memory cost is lower than the adjoint path-kernel method, yet it still computes the full gradient with cost independent of the number of parameters.
@@
    Compared with the finite-element method in [12], its main expression is pathwise, so we can compute it via Monte Carlo algorithms in high dimensions.
@@
    Compared with the likelihood ratio method in [15], the diffusion coefficient $\sigma$ can depend on the parameter $\gamma$ and the location $x$.
@@
    Compared with the fast response method in [33], it does not require hyperbolicity (in the sense of dynamical systems).
@@
    The size of the expression does not grow exponentially to $T$, so we can easily extend our result to stationary distributions (see Section˜3.4).
  6. (6)
    Compared with the author’s divergence-kernel method for scores in [31], this paper considers the parameter-gradients, which enables a new framework of parametric SDE generative models.
This formula gives an efficient gradient tool for diffusion processes with not too small noise or not too strong contraction. In Section˜4.2, the corresponding Monte-Carlo algorithm computes the linear response of the marginal density of a Lorenz 96 system, which is difficult for other methods.
With the new divergence-kernel tool, we develop a new framework for generative models. The logic is rather simple. Our model is an SDE, whose drift and diffusion depends on the state and (many) undetermined parameters. The loss function is the KL-divergence of the empirical distribution of the data with respect to the marginal distribution of the SDE. With the new gradient tool, we can just optimize parameters in the SDE model via stochastic gradient descent.
This new framework combines several advantages of different current generative models, and also offers some new advantages:
  1. (1)
    Directly uses the KL divergence between the empirical data distribution and the marginal density of the SDE as the training objective.
  2. (2)
    Accommodates parametrizations in both drift and diffusion, so prior knowledge of the SDE model can be easily incorporated into the model and reduce the complexity (number of parameters) of the model.
  3. (3)
    Allows deep models (large $T$) without imposing normalizing requirements.
  4. (4)
    It only saves the current state. In particular, we do not have a backward procedure, so we do not save the entire path, substantially reducing the memory cost.
  5. (5)
    If we use the stationary measure of the SDE model in the framework, then the ergodic formula in Section˜3.4 further improves efficiency.
Finally, we might achieve better overall efficiency in general situations; but this requires many more experiments to verify.
@@ 1.2.1. Linear response methods
The path-perturbation method computes the path-perturbation under fixed randomness (such as Brownian motion), then averages over many realizations of randomness. This is also known as the ensemble method or the stochastic gradient method [10, 24]. It also includes the backpropagation method, which is the basic algorithm for machine learning. However, it is expensive for chaotic or unstable systems; the work-around is to artificially reduce the size of the path-perturbation, such as shadowing or clipping methods [34, 37, 39, 11], but they all introduce systematic errors. Moreover, it computes the derivative of an averaged observable, but does not give the derivative of the probability density.
The divergence method is also known as the transfer operator method, since the perturbation of the measure transfer operator is some divergence [14, 46, 48]. However, it is expensive for high-dimensional systems with contracting directions. For SDEs, we can take the derivative of the Fokker-Planck equation, which involves the divergence of the drift and a second-order differential operator [18], which is hard to sample on paths.
The kernel-differentiation method works only for random systems; as the name indicates, the derivative now hits the probability kernel. It is also known as the Cameron-Martin-Girsanov theorem [6, 25], the likelihood ratio method [42, 40, 15], or the Monte-Carlo gradient method [5]. However, it cannot handle multiplicative noise or perturbation on the diffusion coefficients. It is also expensive when the noise is small. We gave an ergodic version and a foliated version in [32].
@@
Our path-kernel method in [30] gives the pathwise linear response formula of the diffusion coefficients. It works well when the noise level is not too small or the largest Lyapunov exponent is not too large; it also does not require hyperbolicity. The adjoint version solves a difficult version of variational data assimilation problem in chaotic systems with unknown parameters and partial observation, where the loss is a single long-time functional accounting for discrepancies in both the observations and the dynamics. However, it can be expensive when the noise is small and the largest Lyapunov exponent is large. Moreover, only part of the expression is derivative on the terminal densities. The path-kernel idea was also used in the Bismut-Elworthy-Li formula for the derivative with respect to the initial conditions [4, 9, 41].
@@
Our divergence-kernel method in [31] gives the score (spatial-derivative of marginal density) of random dynamical systems and SDEs. Compared to previous pathwise expressions for scores [45, 22, 26], the main advantage is that it works for multiplicative noise efficiently. It is good at systems with not too much contraction or not too little noise. However, it can be expensive when the noise is small and the contraction is strong. It fits into the conventional framework of score-based diffusion models, and it enables us to use nonlinear SDEs with multiplicative noise as the forward model, since it can compute the scores efficiently. This paper extends this result to linear responses, which further enables a new framework of generative models.
1.2.2. Generative models
Our new linear response formula opens up the possibilities of redesigning the current framework of diffusion models. At a high level, the bottleneck in ‘parametric SDE as a generative model’ is that the most direct training loss for i.i.d. data is the KL-divergence. But optimizing this loss requires differentiation of the model’s marginal density. For general SDEs this was typically intractable, where many classical tricks either break or become prohibitively high-variance. So previously we had to either replace the training loss or simplify the SDE.
A first workaround is to keep using a parameterized SDE to learn a distribution, but replace KL-divergence training loss by the discrepancy over expectations of certain observables; this covers adversarial critics, kernel MMD, and moment matching [1, 17, 20]. This new loss does not require evaluating the terminal marginal density or its parameter-gradient, so we can use the standard backpropagation to compute the gradient of this loss. However, in chaotic or unstable settings, we need to replace conventional backpropagation method by new methods such as the adjoint path-kernel method [29]. More importantly, the fit of the data is limited by the chosen observable or the class holding the observable: the user needs to introduce an additional modeling choice (the observable/critic/kernel/features/weights) that mediates what aspects of the distribution are matched.
A second workaround, prominent in diffusion models [44] and flow-matching methods [21, 49], is to introduce an auxiliary probability path: a designer-chosen time-indexed family of intermediate probability distributions $\{H_{t}\}_{t\in[0,T]}$ (typically $T=1$) connecting a simple base distribution $H_{0}$ to the data distribution $H_{T}$. The main object to be learned now is the score or a velocity field, and the corresponding training loss becomes the score/denoising loss or vector-field regression loss. Similar to the first workaround, this training loss can be differentiated by standard backpropagation. However, the auxiliary probability path is typically decided without consulting the prior structural knowledge about the model, so the prior knowledge does not automatically translate into computational savings. Even if we have a very simple SDE with a few parameters, the score is not necessarily simple, so the model for the score still needs a large network and potentially many training steps.
A third workaround is to keep using the KL divergence as the loss, but restrict the model class so that the dependence of the distribution on the parameters becomes tractable. We give three examples. First, autoregressive models factorize the joint density as a product of conditional densities under a chosen ordering; this introduces an additional modeling choice (the ordering) and typically requires causal/masked architectural constraints to parameterize these high-dimensional conditionals [38]. Second, in normalizing flows [7] one represents the model distribution through an invertible transport map, so that the change-of-variables formula expresses the density via a Jacobian determinant. Gradients of the resulting log-determinant terms are closely related to divergence-type expressions, connecting to the divergence (transfer-operator) method in linear response (see the review in Section˜1.2.1); in practice, moderately deep flows often require additional constraints or regularization to maintain good conditioning. Third, one may fix the diffusion coefficient, which makes likelihood-based gradients more tractable via likelihood-ratio/Girsanov-type estimators or related score-function identities. However, this narrows the model class: if the diffusion cannot be optimized, the induced SDE marginals may remain far from the data distribution.
Summarizing, our divergence-kernel method enables the gradient tool to optimize KL divergence with respect to the diffusion coefficient or under multiplicative noise. With this, we can work with the simplest generative model, without introducing an auxiliary probability path or an auxiliary observable and without freezing the diffusion. This setting is common rather than exceptional: in mechanistic SDE modeling one often wants likelihood-style marginal fitting while allowing parameters in the diffusion; existing approaches typically avoid this combination.
@@ 1.3. Structure of paper
Section˜5 proposes a parametric SDE generative model, DK-SDE (DK is short for divergence-kernel), which involves only two forward processes. It uses the divergence-kernel formula (not the backpropagation method, which belongs to the path-perturbation method) to perform a gradient descent on the parameters of an SDE model. Its loss function is the KL-divergence between the data and the marginal distribution of the SDE. The logic is more straightforward than the current diffusion models, and the memory cost is less.
@@ 2. Notations and assumptions
where $\left|g_{*}\right|$ is the determinant of the matrix. It is called a divergence due to an equivalent definition by contracting the Hessian (see [35]). Note that this is not the row-wise (or column-wise) divergence of a matrix.
@@
The density of the initial distribution $h_{0}$, the probability kernel $k$, drift $f$, and diffusion $\sigma$ are $C^{1}$ jointly in $\gamma$ and $x$. The diffusion field $\sigma^{\gamma}(x)$ is uniformly bounded away from zero.
Our derivation is rigorous for the discrete-time case, but is formal when passing to the continuous-time limit, so we do not list the technical assumptions here, and leave the rigorous proof to a later paper. We just assume that all integrations, averages, and change of limits are legit. All numerical experiments use the time-discretized formulas; the continuous-time statements serve as guiding limits.
@@ 3.1. Divergence formula One-step system
That being said, we can certainly mix the kernel and divergence linear response formulas in one-step, then repeatedly apply to many-step systems, and then pass to the continuous-time limit. That approach might be beneficial if we later work with anisotropic noise, where $\sigma$ is a matrix, and each parameter controls only a low-dimensional structure, so we do not have the $M$ factor. Another scenario where this approach might be more useful is that the noise is low-dimensional. Both situations incur other difficulties, so we will defer to later papers.
@@ 3.2. Discrete-time divergence-kernel formula for linear responses
$\begin{split}\nu_{0}=\nabla\log h_{0}(x_{0}),\qquad\nu_{n+1}=(1-\alpha_{n+1})g_{b_{n}}^{*-1}\left(\nu_{n}-\textnormal{div}g_{b_{n}*}(x_{n})\right)+\frac{\alpha_{n+1}\nabla\log k(b_{n})}{\sigma_{n}(x_{n})}.\end{split}$
@@
Note that we need the score of $h_{0}$, or $\nu_{0}$. Only for large $T$ and systems with fast decorrelation, we can use any $\nu_{0}$ and the impact would be negligible. See Section˜3.4 for more details.
@@ 3.3. Preparations for SDEs
We keep only terms of order $\Delta t$} in expectation and drop higher-order terms since they integrate to zero; this is standard in Ito calculus. Ignoring terms that go to zero in the continuous-time limit, [31] derived explicit expressions of some terms for time-discretized SDEs,
@@
[…]
For the SDE in $\mathbb{R}^{M}$,
@@ 3.4. Continuous-time divergence-kernel formula for linear response
Then substitute to derive the formula. ∎
[…]
Ignoring high-order terms which vanish as ${\Delta t}\rightarrow 0$,
@@
With these, we can formally pass Definition˜1 to the continuous-time limit to derive the main formula of this paper.
Derivation.
(of Section˜1.1) Substitute Sections˜3.4, 3.4 and 3.4 into Definition˜1, we get
@@
We also formally pass to the infinite-time limit of stationary distributions $h^{\gamma}$. We need to further assume that $F_{t}$ and $\sigma_{t}$ are stationary, and that there is $T^{\prime}$ such that $\alpha_{t}$ only depends on $x_{t}\sim x_{T^{\prime}}$, so $(\alpha,\nu,x)$ are jointly stationary.
[…]
For any $C^{2}$ observable function $\Phi:\mathbb{R}^{M}\rightarrow R$, let $h^{\gamma}$ be the stationary measure under parameter $\gamma$, then the linear response of the averaged observable can be expressed almost surely on one path:
@@ 3.5. Discussions
The divergence-kernel method contains second derivatives. To speed up our code, note that we do not really need the matrices such as $\nabla^{2}\sigma$ and $\nabla F^{T}$, we only need their products with certain vectors. Such matrix-vector products can be implemented much more efficiently in high dimensions (such as the jvp function in JAX) than forming the matrix, especially with sparse structures.
@@
4. Numerical examples for linear responses
@@
$\begin{split}dx=\beta_{t}(\gamma-x)dt+\sqrt{\beta_{t}}\left(0.5+\exp\left((\gamma-x)^{2}\right)\right)dB,\quad\textnormal{where}\quad\beta_{t}:=1+3t.\end{split}$
The initial distribution is $x_{0}\sim N(0,1)$. Set $T=1$, ${\Delta t}=0.01$, $\alpha=10$. Note that the problem is symmetric for $\gamma$.
We are only interested in terminal states $x_{T}$ in the interval [2,2][-2,2], which is further partitioned into $10$ subintervals. We simulate many sample paths, then count the number of paths whose $x_{T}$ fall into a specific subinterval, to obtain the empirical distribution of $x_{T}$ on subintervals. Then we compute the linear response $\delta\log h_{T}$ by averaging the divergence-kernel formula over these paths, and see if our results reflect the trend of $\log h_{T}$ from the empirical distribution. We also use the average of $\nu_{T}$ to compute the scores $\nabla\log h_{T}$. Each subinterval contains at least 180 samples. The $\delta\log h_{T}$ and $\nabla\log h_{T}$ computed for different $\gamma$ and $x$ are shown in Figure˜1. Both correctly reflect the trend of $\log h_{T}$.
Figure 1. Divergence-kernel method for the linear responses and scores of the 1-dimensional SDE with multiplicative noise, $T=1$. The dots are $\log h_{T}$. Each short line is a linear response or score computed by the divergence-kernel algorithm, averaged on paths whose terminal states $x_{T}$ fall into the same subinterval. Left: linear responses $\delta\log h_{T}$. Right: scores $\nabla\log h_{T}$ .
@@
We use Section˜3.4 to compute the linear response of the stationary density of the Lorenz 96 model [23]. The dimension of the system is $M=40$. The SDE is
$dx^{i}=\left(\left(x^{i+1}-x^{i-2}\right)x^{i-1}-x^{i}+8+\gamma-0.01(x^{i})^{2}\right)dt+\left(0.5+\exp\left(|x-\gamma|^{2}\right)\right)\,dB^{i},$
$\textnormal{where}\quad i=0,\ldots,M-1;\quad x_{0}\sim{\mathcal{N}}(0,I).$
Here $i$ labels different directions in $\mathbb{R}^{M}$, and We use cyclic indexing modulo $M$. We added noise and the $0.01(x^{i})^{2}$ term, which prevents the noise from carrying us to infinitely far away. We choose the observable function to be
@@
We use the Euler integration scheme with ${\Delta t}=0.002$. A typical orbit is in Figure˜2. Then we use the ergodic linear response formula Section˜3.4 to compute the linear response of the observable averaged on the stationary measure. The averaged observable $\Phi_{avg}$ and the linear response $\delta\int\Phi h^{\gamma}$ are averaged over 7 orbits of length $T=400$. The decorrelation time is $W=1.5$. The results shown in Figure˜2; the linear response computed using the divergence-kernel method is correct.
@@
5. Parametric SDE generative model trained by the divergence-kernel method (DK-SDE)
We consider the problem in which we are given $N_{\mathrm{data}}$ many samples $y_{k}\in\mathbb{R}^{M}$, whose true distribution is unknown. Then we use the divergence-kernel method to compute the gradient, and use gradient descent to train a parametric generative SDE model which maps the Gaussian distribution to match the data.
5.1. Problem and algorithm description
We describe the problem and our model, then we define loss function and explain how to compute its gradient, and how to update the parameters by gradient descent.
First, we describe how to generate samples. Note that the information in this paragraph will not be available when we train the model. The samples are generated by the SDE